A person scuffing her feet on a wool rug on a dry day accumulates a net charge of -89?C .


1.) How many excess electrons does she get?


2.) By how much does her mass increase?
Answer #1

Given: Charge accumulated q-89uC Charge accumulated- No. of e gained*charge of electron 1. -89*10%N,*-1.6*10-19 Ne -5.5625 10
Answer #2

Charge = -89x10^-6 C

Charge/e = -1.6x10^-19 C

No. of electrons = 89x10^-6 / 1.6x10^-19 = 55.625x10^13 electrons


Mass/e = 9.11x10^-31 kg

Mass of electrons = 9.11x10^-31 * 55.625x10^13 = 506.74x10^-18 kg
Answer #3

The charge on the electron e = -1.602 x 10-19 C. The person in your problem has accumulated -89 x 10-6 C. The number of electrons required to produce this charge is

-89 x 10-6 C / -1.602 x 10-19 C / electron =55.625 x 10^13 electrons


The mass of the electron is 9.11 x 10-31 kg. Multiply the number you just got by that mass to find the increase in mass


=5.067x10^-16kg
Answer #4

excess electrons = -89*10^-3/-1.6*10^-19=55.625*10^16 elECtrons

b)MASS INCREASED = 9.1*10^-31*55.625*10^16=50.6*10^-14 kg
Answer #5

Nel = 89e-6 C/ 1.6e-19 C= 55625 electrons

mel = 55625 X 9.1e-31 kG =5.06x10^-20kg
Answer #6

\(-89\mu C * \frac{1*10^{-6}C}{1\ \mu C}*\frac{1 e^{-}}{-1.6*10^{-19}C} = 556,250,000,000,000 electrons \)   

    \(Mass = 556,250,000,000,000 *(9.11*10^{-31}kg) = 5.067*10^{-16}kg\)   
Answer #7

Q = -60